OpenStudy (anonymous):

someone help me solve this integrate k(1+k^2)^-3

OpenStudy (anonymous):

Do you want to integrate this or what is the question?

OpenStudy (anonymous):

yeah u integrate

OpenStudy (anonymous):

\[\int\limits\frac{k}{(1+k^{2})^{3}} \]?

OpenStudy (anonymous):

aren't we missing a dk here?

OpenStudy (anonymous):

Yes, lol \[\int\limits\frac{k}{(1+k^{2})^{3}} dk\] But need the asker you confirm if the integral is right.

OpenStudy (anonymous):

Anyway, so what you need to do is a U substitution. Have you done this before?

OpenStudy (anonymous):

Integration by substitution I mean.

OpenStudy (anonymous):

it should be integration by parts

OpenStudy (anonymous):

Doesn't look like it needs to be. You have two algebraic functions. If you substitute twice, you can get an answer, unless they specifically asked for the solution using integration by parts?